OpenStudy (anonymous):

function f & g are both concave fns of a single variable. Neither fn is necessarily differentiable. is the fn defined by h(x)=f(x)+g(x) necessarily concave, necessarily convex or not necessarily either.

OpenStudy (anonymous):

I am thinking it would be "necessarily concave"... what do you think? I am still considering what impact comes from the fact that it says the functions f and g are not necessarily differentiable.

OpenStudy (anonymous):

that wer my doubt area is.. cz wen u jst use the fact that f n g are concave u can show h is concave but wat difference does differentiability makes here

OpenStudy (anonymous):

yes... interesting. What would make f not differentiable but still allow it to be considered concave? There is the "easy" definition of concave as U shaped (bowl, open side up). But the definition based on derivatives depends, I would have thought, on the function being differentiable. Maybe you can imagine a non-differentiable function that still faces upward.... if you can imagine one, then a second similar upward facing function probably (but I hate guessing!!) can just be added to the first one without affecting concavity.

OpenStudy (anonymous):

certainly if you had f(x) = x^2 and g(x) = 2x^2, adding them to get h(x) = 3x^2 is still concave. (wait, am I totally backwards on concave? an upward parabola is concave up, right? Not convex? It's been awhile since I've done this sort of problem).

OpenStudy (anonymous):

ya upward parabola is concave up

OpenStudy (anonymous):

concave fns hav minima ie second differenciation is >0

OpenStudy (anonymous):

so, f(x) = x^2 is concave up, but it doesn't demonstrate non-differentiability. But I brought it up as a simpler example... adding 2 concave-up parabolas results in a 3rd up-facing parabola

OpenStudy (anonymous):

v need to prove tis ... how shud i go abt it

OpenStudy (anonymous):

so, leaving aside the non-differentiable part for a sec, I'm pretty certain you could prove analytically that for any f(x) and g(x) that are concave up, h(x) is also concave up.

OpenStudy (anonymous):

ya.

OpenStudy (anonymous):

do you actually have to prove this? Or just answer? Also, help me on this non-differentiable idea... what makes something non-differentiable? Is it that a derivative is undefined, like a section of vertical slope?

OpenStudy (anonymous):